OpenStudy (anonymous):

4/sqrt20 simplify please help

OpenStudy (callisto):

First, simplify sqrt20 \[\sqrt{20} = \sqrt{4\times 5} =\sqrt{2^2\times 5} = ...?\]

OpenStudy (anonymous):

√20/5 this is what I got?

OpenStudy (callisto):

Ah... Not the simplified one, I think... Can you simplify sqrt20?

OpenStudy (anonymous):

Yes and no

OpenStudy (anonymous):

Lol well 5 goes into 20 Cz 4x5 and 5x1 so wold it be sqrt4/1 ???

OpenStudy (callisto):

Not really... \[\frac{4}{\sqrt{20}} = \frac{4\sqrt{20}}{(\sqrt{20})^2}=\frac{4\sqrt{20}}{20}=\frac{\sqrt{20}}{5} = \frac{\sqrt{4\times 5}}{5}=\frac{\sqrt{2^2\times 5}}{5} = \frac{\sqrt{2^2}\times \sqrt{5}}{5} = ...?\]

OpenStudy (krabsen):

Callisto can i ask you about something about this question when you are done

OpenStudy (callisto):

Yes?!

OpenStudy (anonymous):

Wouldn't the sqrt cancel out the exponent?

OpenStudy (anonymous):

can i ask other questions while your answering?

OpenStudy (callisto):

@22makayla22 Hmm... when exponent = 2 and there this a sqrt, you can.

OpenStudy (callisto):

*there is

OpenStudy (anonymous):

hmmm very confusing

OpenStudy (callisto):

For example \[\sqrt{3^2} = 3\]

OpenStudy (anonymous):

so then would it just be sqrt2?

OpenStudy (anonymous):

or 2?

OpenStudy (callisto):

What do you think? @22makayla22

OpenStudy (anonymous):

i guess 2?

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

Yes. So, what do you get for the answer?

OpenStudy (callisto):

Yes for \(\sqrt{2^2}\) = 2 :|

OpenStudy (richyw):

The easiest way to do this is just write the numerator as the square root of its square. Then divide what's inside the square root and simplify from there\[=\sqrt{\frac{16}{20}}=\sqrt{\frac{4}{5}}= \frac{2}{\sqrt{5}}\]

OpenStudy (richyw):

well I mean the easiest way is just to pull out a 4

OpenStudy (richyw):

either way it should take you two steps

OpenStudy (anonymous):

@Callisto 2/srt5 because the top one cancels

OpenStudy (anonymous):

@richyw can i simplify it more or is that it?

OpenStudy (callisto):

You don't have to rationalise the answer? @22makayla22

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

thhank you so much @richyw and @Callisto

OpenStudy (krabsen):

4/sprt20= 4/(20^1/2)= 4*1/(20^1/2)= 4*20^-1/2 cant you do it like that?

OpenStudy (anonymous):

and @Krabsen ?? thank you

OpenStudy (krabsen):

hehe just another way to do it

OpenStudy (richyw):

well you could multiply the top and bottom by \(\sqrt{5}\)\[\frac{2 \sqrt{5}}{5}\]

OpenStudy (anonymous):

oh okay i get it thanks @Krabsen

OpenStudy (anonymous):

THANK YOU EVERY ONE

OpenStudy (callisto):

I'm not sure.. But when I have to simplify something like that, I have to do the rationalisation too. That is: \[\frac{4}{\sqrt{20}} = \frac{4}{2\sqrt{5}} =\frac{2\sqrt5}{5}\]